if cosecA + cotA = 11/2, then tanA =?
a)21/22               b)15/16
c)44/117             d)117/43

Grade:11

1 Answers

Divyanshi
6 years ago
cosecA+cotA=1/sinA+cosA/sinA =(1+cosA)/sinA 11/2=2(cosA/2cosA/2)/(2sinA/2.cosA/2)11/2=(cosA/2)/(sinA/2)11/2=cotA/22/11=tanA/2Now.... tanA=(2tanA/2)/1-tanA/2tanA/2 tanA=(2*2/11)/(1-4/121) tanA=(4/11)/(117/121) tanA=(4*11)/117 tanA=44/117